Please help me. I have tried, but have not yet succeded! Here's the problem: The sum of three consecutive even integers is -24. Find the largest integer.
Answer by atif.muhammad(135)

The sum of three consecutive even integers is -24. Find the largest integer.
Let the first integer be x, the next consecutive even integer be, (x+2) and the third consecutive even integer be (x+4).
Their sum = -24
x + (x+2) + (x+4) = -24
3x + 6 = -24
3x = -30
x = -10
(x+2) = -8
(x+4) = -6
The largest integer is -6.